# Spline Dimensional Decomposition with Interpolation-based Optimal Knot Selection for Stochastic Dynamic Analysis

### v1

@@ 4.1 Example 1: frequency response analysis of a two-degree-of-freedom system
Two distinct cases, one considering the randomness of the spring stiffness only $(N=1)$ and other allowing the randomness of mass, damping, and stiffness properties $(N=3)$, are studied. This example aims to perform uncertainty quantification of the displacement amplitude $|Z_{1}(\omega;\mathbf{X})|$ in the frequency domain using both the SDD and GP methods. First, standard deviation is estimated over a range of excitation frequencies to capture the effect of input uncertainty. Then, the failure probability is calculated using $10^{6}$ Monte Carlo resamples applied to the constructed SDD and GP surrogate models.
@@
$p=2$ $I=8$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.25,-1.5,\\ -0.75,0,0.75,1.5,2.25,3,\\ 3,3\}\end{array}$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.25,-1.25,\\ 0.06,0.3,1.25,2.36,2.28,\\ 3,3,3\}\end{array}$
$I=16$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.625,-2.25,\\ -1.875,-1.5,-1.125,-0.75,\\ -0.375,0,0.375,0.75,1.125,\\ 1.5,1.875,2.25,2.625,3,\\ 3,3\}\end{array}$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.94,-2.77,\\ -2.5,-2.2,-2.15,-2.0,-1.69,\\ -1.51,-0.93,-0.49,-0.03,0.01,\\ 1.19,2.27,2.71,3,3,3\}\end{array}$
@@ 4.1.1 Case 1: Randomness in stiffness ($N$=1)
The system parameters are defined as $M_{1}=M_{2}=1\ [\mathrm{kg}]$, $C_{1}=C_{2}=1\ [\mathrm{N}/\mathrm{ms}]$, $K_{1}=K_{2}=15,000(1+X)\ [\mathrm{N}/\mathrm{m}]$, where the random variable $X$ follows a uniform distribution over $[-3,3]$. To assess the accuracy of the SDD approximation, univariate models with $p=1,2$ and $I=8,16$ are constructed using three type of knot vector: uniformly spaced, randomly spaced, and the proposed interpolation-based method. Table 1 summarizes the corresponding knot configurations. Once the models are built, the standard deviation of displacement $z_{1}(\omega;X)$ is computed and compared against the reference solution obtained from $10^{5}$ crude MCS simulations. Table 2 shows the mean errors of the standard deviation over all frequencies.
@@
Figure 5: Standard deviation of $|Z_{1}(\omega;X)|$ over the frequency range from 10 [Hz] to 35 [Hz] for uniformly spaced, randomly spaced, and proposed SDD configurations compared against 100,000 Monte Carlo simulations. The results are presented for (left) $p=1$ and $I=16$, and (right) $p=2$ and $I=16$. (Case 1)
@@
The second natural frequency corresponding to the mean input parameters is 31.465 [Hz]. To further assess dynamic behavior, cumulative distribution functions (CDFs) of $|Z_{1}(\omega;X)|$ at this frequency are evaluated using $10^{6}$ MCS samples drawn from the constructed surrogates. Figure 6 presents that the SDD model with the proposed knot vector accurately captures the response distribution, closely matching the reference MCS results. However, in this low-dimensional case, the GP achieves slightly better predictive accuracy than the SDD methods.
Figure 6: Cumulative probability distribution functions of $|Z_{1}(\omega;X)|$ at 31.465 [Hz], when $p=2$ and $I=16$. The results are presented for (left) displacement $|Z_{1}(\omega;X)|$ ranges 0 to 15, and (right) ranges 5.6 to 14.4. (Case 1)
We assess the probability of failure at the second natural frequency (31.465 [Hz]) with a displacement threshold of $10^{-4}$. When any realization of $|Z_{1}(\omega;X)|$ exceeds this threshold, we consider it a failure. Table 3 presents the failure probabilities computed using different surrogate models, varying the polynomial order $p$, number of internal knots $I$, and the knot vector type (uniform, random, or the proposed optimal method). For comparison, the reference probability from crude MCS is also shown.
In all cases, the proposed SDD method demonstrates better agreement with the reference failure probability from crude MCS than both uniform and random knot vector configurations. In particular, the proposed method shows reduced deviation from the MCS-based benchmark, validating its robustness in failure probability estimation. Although GP presents slightly higher accuracy than SDD, optimal knot selection consistently improves predictive accuracy within the SDD framework.
Table 3: Comparison of failure probabilities $\mathbb{P}[|Z_{1}(\omega;X)|>10^{-4}]$ estimated using SDD with different internal knot configurations and GP model at $\omega=31.465$ [Hz]. Results are provided for various polynomial orders ($p$) and numbers of internal knots ($I$), with crude MCS results as reference (Case 1).
@@ 4.1.2 Case 2: Randomness in mass, damping, and stiffness (N=3)
We define the parameters $M_{1}=M_{2}=(1+0.05X_{M})\ [\mathrm{kg}]$, $C_{1}=C_{2}=(1+0.05X_{C})\ [\mathrm{N}/\mathrm{ms}]=$, and $K_{1}=K_{2}=15,000(1+0.05X_{K})[\ \mathrm{N}/\mathrm{m}]$, where random variables $\mathbf{X}=(X_{M},X_{C},X_{K})^{\intercal}$ follow a truncated Gaussian distribution bounded in $[-3,3]$ with standard deviations $(\delta_{M}=\delta_{C}=\delta_{K}=0.05)$. Bivariate SDD models with $p=1,2$ and $I=8$ are constructed using three different knot configurations: uniformly spaced, randomly spaced, and the proposed optimal knots. For both uniformly and randomly spaced knots, an identical knot vector is applied across all directions $k=1,2,3$, after which the standard deviation of $|Z_{1}(\omega;\mathbf{X})|$ is evaluated.
Figure 7: Standard deviation of $|Z_{1}(\omega;\mathbf{X})|$ over the frequency range from 10 [Hz] to 35 [Hz] for uniformly spaced, randomly spaced, and proposed SDD configurations ($p=2$, $I=8$) compared against 100,000 Monte Carlo simulations. The results are presented for frequencies (left) 10 to 35 [Hz], and (right) 10.9 to 13.9 [Hz]. (Case 2)
Figure 7 demonstrates that the SDD model with the proposed knots, using $p=2$ and $I=8$, shows excellent agreement with the 100,000 sample MCS reference throughout the entire frequency range. In the 10.9–13.9 [Hz] region, where rapid variations occur, the proposed method accurately captures the nonlinear trend, while the uniformly spaced SDD tends to overestimate the standard deviation.
@@
Figure 8 presents the cumulative distribution functions (CDF) of $|Z_{1}(\omega;\mathbf{X})|$ at the second natural frequency (31.465 [Hz]), computed using $10^{6}$ resamples from each surrogate. The failure probability is defined as the probability of the response that exceeds a threshold value of $10^{-4}$. Table 5 compares the failure probabilities computed by each method against the reference values from crude MCS. The proposed SDD model demonstrates excellent agreement with the MCS benchmark, especially when compared to randomly spaced SDD. This confirms its effectiveness in capturing rare-event behavior in complex uncertainty conditions.
@@
Table 5: Comparison of failure probability estimates for $|Z_{1}(\omega;\mathbf{X})|$ at 31.465 [Hz] across different knot configurations and GP model, benchmarked against crude MCS results (Case 2). Failure occurs when displacement exceeds $10^{-4}$. Results are reported for polynomial orders $p=1,2$ and internal knots $I=8,16$.
@@ 4.2 Example 2: Modal analysis of a turbo fan jet engine (N=5)
Figure 9: The 3D model of turbofan jet engine with detailed views highlighting the components and material composition (a) front view and (b) back view (Example 2); A CAD model of the turbofan jet engine; (c) front view of the engine with four input random variables and (d) back view of the engine with single input random variables; An FEA of the engine; (e) a mesh consisting of 71,406 nodes and 29,995 elements in the nominal model (f) 5000 rpm rotational velocity and boundary conditions, including fixed support and cylindrical support.
@@ 4.3 Example 3: Modal analysis of lower control arm (N=10)
Figure 13: An FEA of the lower control arm with bushings (Example 3); (a) loading and boundary conditions, including axial displacement conditions along the x- and z-axes, compression-only support, and force replaced by displacement constraints (x-axis: 1 [mm], y-axis: 2 [mm]); (b) a tetrahedral mesh (the lower control arm) and a hexahedral mesh (the bushing) consisting of 154,855 nodes and 41,286 elements in the nominal model.
@@
To further assess model accuracy, one million input samples were generated based on the ten random variables, and the trained surrogate models were evaluated on this larger set. As illustrated in Figure 14, the SDD model with the proposed knots closely approximated the distribution of both the first and second natural frequencies about MCS. The $R^{2}$ value for the first mode reached 0.984, indicating a 7.7% improvement over the randomly spaced knot model (0.914) and a 2.8% improvement over the uniformly spaced knot model (0.957). A similar performance gain was observed for the second mode. These results confirm that the proposed knot method not only improves variance accuracy but also preserves distributional characteristics with fewer training samples (401 FEA results), compared to the 2,000 MCS results. The number of basis functions and the size of the training dataset are fixed for all knot configurations. As a result, the computational cost required to train the surrogate models remains unchanged. Therefore, the improved accuracy achieved by the proposed method stems solely from the more effective placement of internal knots, not from increased data or model complexity. This highlights the SDD with the proposed method as a statistically efficient and scalable approach for high-dimensional UQ in structural dynamics.
@@ 5 Conclusions
This study proposed a novel optimal spline dimensional decomposition (SDD) method that incorporates an optimal knot vector for uncertainty quantification in dynamic systems. While B-spline basis functions are effective for modeling non-smooth responses, their performance is highly sensitive to the placement of internal knots. To address this issue, we developed an interpolation-based knot selection method that avoids additional optimization. The method first approximates the input-output data profile using linear interpolation. It then identifies reference regions exhibiting high gradient variation, which are particularly important in high-dimensional problems with oscillatory responses. Within each region, the point of maximum gradient is selected as an internal knot, thereby constructing an optimal knot vector tailored to the stochastic characteristics of the response. This optimal knot vector is then embedded into the spline basis to develop an optimal SDD model.
The effectiveness of the proposed method is validated through three numerical examples, which include frequency and modal analyses with up to ten random variables. Compared to uniformly and randomly spaced knots, the proposed approach consistently yielded lower variance errors, and estimated cumulative distribution functions (CDFs) are in good agreement with MCS. The proposed method proved its effectiveness in solving higher dimensional scenarios that often apply to industrial structures such as a turbofan jet engine and lower control arm. While the proposed SDD model showed slightly lower accuracy than Gaussian Process (GP) models in low-dimensional cases, it outperformed GP in the case of increased input dimension. These results demonstrate the method’s capability to capture strong non-linearities using appropriately placed knots, even with limited data. Consequently, the method successfully estimated failure probability estimates that are in close agreement with MCS.
However, the current approach may limit accuracy for uncertain response shapes as it estimates input-output profiles using linear interpolation. In addition, the method cannot currently predict frequency-domain responses in a unified manner. Incorporating higher-order interpolation schemes (e.g., quadratic or cubic) can improve accuracy. Future work will address these limitations by exploring advanced interpolation methods and integrating intrusive dynamic system modifications to enable full-spectrum frequency predictions. Ultimately, the proposed method’ potential in design optimization will be further evaluated by embedding reliability indices within RBDO frameworks.
@@ B.1 Ogden model and Prony series
Ogden model is [14]
$U=\sum_{i=1}^{N}\frac{2\delta_{i}}{\alpha_{i}^{2}}\left(\lambda_{1}^{\alpha_{i}}+\lambda_{2}^{\alpha_{i}}+\lambda_{3}^{\alpha_{i}}+\lambda_{4}^{\alpha_{i}}-3\right)+\sum_{i=1}^{N}\frac{1}{D_{i}}\left(U-1\right)^{2},$ (36)
where $\bar{\lambda}_{n}$ represents the principal stretches, $J$ is the volume ratio, $\delta_{i}$ and $\alpha_{i}$ are determined by a fitting with the stress-strain curve of the material, and $D_{i}$ means the compressibility. Prony series for viscoelastic is  [14]
$G(t)=G_{0}\left[1-\sum_{i=1}^{N}g_{i}e^{\left(\frac{-t}{\tau_{i}}\right)}\right],$ (37)
where $G(t)$ and $G_{0}$ are the shear modulus and the initial shear modulus. The terms $g_{i}$ and $\tau_{i}$ represent the dimensionless shear-relaxation modulus and the relaxation time.

### v2

@@ 4.1 Example 1: frequency response analysis of a two-degree-of-freedom system
Two distinct cases, one considering the randomness of the spring stiffness only $(N=1)$ and other allowing the randomness of mass, damping, and stiffness properties $(N=3)$, are studied. This example aims to perform uncertainty quantification of the displacement amplitude $|Z_{1}(\omega;\mathbf{X})|$ in the frequency domain using both the SDD and GP model with a Matern kernel ($l=1.0$, $\nu=0.5$). First, standard deviation is estimated over a range of excitation frequencies to capture the effect of input uncertainty. Then, the failure probability is calculated using $10^{6}$ Monte Carlo resamples applied to the constructed SDD and GP surrogate models.
@@
$p=2$ $I=8$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.25,-1.5,\\ -0.75,0,0.75,1.5,2.25,3,3,3\}\end{array}$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.25,-1.25,\\ 0.06,0.3,1.25,2.36,2.28,3,3,3\}\end{array}$
$I=16$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.625,-2.25,\\ -1.875,-1.5,-1.125,-0.75,\\ -0.375,0,0.375,0.75,1.125,\\ 1.5,1.875,2.25,2.625,3,3,3\}\end{array}$ $\begin{array}[]{l}\xi=\{-3,-3,-3,-2.94,-2.77,\\ -2.5,-2.2,-2.15,-2.0,-1.69,\\ -1.51,-0.93,-0.49,-0.03,0.01,\\ 1.19,2.27,2.71,3,3,3\}\end{array}$
@@ 4.1.1 Case 1: Randomness in stiffness ($N$=1)
The system parameters are defined as $M_{1}=M_{2}=1\ \mathrm{kg}$, $C_{1}=C_{2}=1\leavevmode\nobreak\ \mathrm{N}/\mathrm{ms}$, $K_{1}=K_{2}=15,000(1+X)\ \mathrm{N}/\mathrm{m}$, where the random variable $X$ follows a uniform distribution over $[-3,3]$. To assess the accuracy of the SDD approximation, univariate models with $p=1,2$ and $I=8,16$ are constructed using three types of knot vector: uniformly spaced, randomly spaced, and the proposed interpolation-based method. Table 1 summarizes the corresponding knot configurations. Once the models are built, the standard deviation of displacement $z_{1}(\omega;X)$ is computed and compared against the reference solution obtained from $10^{5}$ crude MCS simulations. Table 2 shows the mean errors of the standard deviation over all frequencies.
@@
Figure 5: Standard deviation of $|Z_{1}(\omega;X)|$ over the frequency range from 10 Hz to 35 Hz for uniformly spaced, randomly spaced, and proposed SDD configurations compared against 100,000 Monte Carlo simulations. The results are presented for (left) $p=1$ and $I=16$, and (right) $p=2$ and $I=16$. (Case 1)
@@
The second natural frequency corresponding to the mean input parameters is 31.465 Hz. To further assess dynamic behavior, cumulative distribution functions (CDFs) of $|Z_{1}(\omega;X)|$ at this frequency are evaluated using $10^{6}$ MCS samples drawn from the constructed surrogates. Figure 6 shows that the SDD model with the proposed knot vector accurately captures the response distribution, closely matching the reference MCS results. However, in this low-dimensional case, the GP achieves slightly better predictive accuracy than the SDD methods.
Figure 6: Cumulative probability distribution functions of $|Z_{1}(\omega;X)|$ at 31.465  Hz, when $p=2$ and $I=16$. The results are presented for (left) displacement $|Z_{1}(\omega;X)|$ ranges 0 to 15, and (right) ranges 5.6 to 14.4. (Case 1)
We assess the probability of failure at the second natural frequency (31.465 Hz) with a displacement threshold of $10^{-4}$. When any realization of $|Z_{1}(\omega;X)|$ exceeds this threshold, we consider it a failure. Table 3 presents the failure probabilities computed using different surrogate models, varying the polynomial order $p$, number of internal knots $I$, and the knot vector type (uniform, random, or the proposed optimal method). For comparison, the reference probability from crude MCS is also shown.
In all cases, the proposed SDD method demonstrates better agreement with the reference failure probability from crude MCS than both uniform and random knot vector configurations. In particular, the proposed method shows reduced deviation from the MCS-based benchmark, validating its robustness in failure probability estimation. Although GP achieves slightly higher accuracy than SDD, optimal knot selection consistently improves predictive accuracy within the SDD framework.
Table 3: Comparison of failure probabilities $\mathbb{P}[|Z_{1}(\omega;X)|>10^{-4}]$ estimated using SDD with different internal knot configurations and GP model at $\omega=31.465$ Hz. Results are provided for various polynomial orders ($p$) and numbers of internal knots ($I$), with crude MCS results as reference (Case 1).
@@ 4.1.2 Case 2: Randomness in mass, damping, and stiffness (N=3)
We define the parameters $M_{1}=M_{2}=(1+0.05X_{M})\ \mathrm{kg}$, $C_{1}=C_{2}=(1+0.05X_{C})\ \mathrm{N}/\mathrm{ms}=$, and $K_{1}=K_{2}=15,000(1+0.05X_{K})\ \mathrm{N}/\mathrm{m}$, where random variables $\mathbf{X}=(X_{M},X_{C},X_{K})^{\intercal}$ follow a truncated Gaussian distribution bounded in $[-3,3]$ with standard deviations $(\delta_{M}=\delta_{C}=\delta_{K}=0.05)$. Bivariate SDD models with $p=1,2$ and $I=8$ are constructed using three different knot configurations: uniformly spaced, randomly spaced, and the proposed optimal knots. For both uniformly and randomly spaced knots, an identical knot vector is applied across all directions $k=1,2,3$, after which the standard deviation of $|Z_{1}(\omega;\mathbf{X})|$ is evaluated.
Figure 7: Standard deviation of $|Z_{1}(\omega;\mathbf{X})|$ over the frequency range from 10 Hz to 35 Hz for uniformly spaced, randomly spaced, and proposed SDD configurations ($p=2$, $I=8$) compared against 100,000 Monte Carlo simulations. The results are presented for frequencies (left) 10 to 35 Hz, and (right) 10.9 to 13.9 Hz. (Case 2)
Figure 7 demonstrates that the SDD model with the proposed knots, using $p=2$ and $I=8$, shows excellent agreement with the 100,000 sample MCS reference throughout the entire frequency range. In the 10.9–13.9 Hz region, where rapid variations occur, the proposed method accurately captures the nonlinear trend, while the uniformly spaced SDD tends to overestimate the standard deviation.
@@
Figure 8 presents the cumulative distribution functions (CDF) of $|Z_{1}(\omega;\mathbf{X})|$ at the second natural frequency (31.465 Hz), computed using $10^{6}$ resamples from each surrogate. The failure probability is defined as the probability of the response that exceeds a threshold value of $10^{-4}$. Table 5 compares the failure probabilities computed by each method against the reference values from crude MCS. The proposed SDD model demonstrates excellent agreement with the MCS benchmark, especially when compared to randomly spaced SDD. This confirms its effectiveness in capturing rare-event behavior in complex uncertainty conditions.
@@
Table 5: Comparison of failure probability estimates for $|Z_{1}(\omega;\mathbf{X})|$ at 31.465 Hz across different knot configurations and GP model, benchmarked against crude MCS results (Case 2). Failure occurs when displacement exceeds $10^{-4}$. Results are reported for polynomial orders $p=1,2$ and internal knots $I=8,16$.
@@ 4.2 Example 2: Modal analysis of a turbo fan jet engine (N=5)
Figure 9: The 3D model of turbofan jet engine with detailed views highlighting the components and material composition (a) front view and (b) back view (Example 2); A CAD model of the turbofan jet engine; (c) front view of the engine with four input random variables and (d) back view of the engine with single input random variables; An finite element analysis of the engine; (e) a mesh consisting of 71,406 nodes and 29,995 elements in the nominal model (f) 5000 rpm rotational velocity and boundary conditions, including fixed support and cylindrical support.
@@ 4.3 Example 3: Modal analysis of lower control arm (N=10)
Figure 13: An finite element model of the lower control arm with bushings (Example 3); (a) loading and boundary conditions, including axial displacement conditions along the x- and z-axes, compression-only support, and force replaced by displacement constraints (x-axis: 1 [mm], y-axis: 2 [mm]); (b) a tetrahedral mesh (the lower control arm) and a hexahedral mesh (the bushing) consisting of 154,855 nodes and 41,286 elements in the nominal model.
@@
To further assess model accuracy, one million input samples were generated based on the ten random variables, and the trained surrogate models were evaluated on this larger set. As illustrated in Figure 14, the SDD model with the proposed knots closely approximated the distribution of both the first and second natural frequencies about MCS. The $R^{2}$ value for the first mode reached 0.984, indicating a 7.7% improvement over the randomly spaced knot model (0.914) and a 2.8% improvement over the uniformly spaced knot model (0.957). A similar performance gain was observed for the second mode. These results confirm that the proposed knot method not only improves variance accuracy but also preserves distributional characteristics with fewer training samples (401 simulation results), compared to the 2,000 MCS results. The number of basis functions and the size of the training dataset are fixed for all knot configurations. As a result, the computational cost required to train the surrogate models remains unchanged. Therefore, the improved accuracy achieved by the proposed method stems solely from the more effective placement of internal knots, not from increased data or model complexity. This highlights the SDD with the proposed method as a statistically efficient and scalable approach for high-dimensional UQ in structural dynamics.
@@ 5 Conclusions
This study proposed a novel optimal spline dimensional decomposition (SDD) method that incorporates an optimal knot vector for uncertainty quantification in dynamical systems. While B-spline basis functions are effective for modeling non-smooth responses, their performance is highly sensitive to the placement of internal knots. To address this issue, we developed an interpolation-based knot selection method that avoids additional optimization. The method first approximates the input-output data profile using linear interpolation. It then identifies reference regions that exhibit high gradient variation, which are particularly important in high-dimensional problems with oscillatory responses. Within each region, the point of maximum gradient is selected as an internal knot, thereby constructing an optimal knot vector tailored to the stochastic characteristics of the response. This optimal knot vector is then embedded into the spline basis to develop an optimal SDD model.
The effectiveness of the proposed method is validated through three numerical examples, which include frequency and modal analyses with up to ten random variables. Compared to uniformly and randomly spaced knots, the proposed approach consistently yielded lower variance errors, and estimated cumulative distribution functions (CDFs) are in good agreement with MCS. The proposed method proved its effectiveness in solving higher-dimensional scenarios that often apply to industrial structures such as a turbofan jet engine and lower control arm. While the proposed SDD model showed slightly lower accuracy than Gaussian Process (GP) models in low-dimensional cases, it outperformed GP in the case of increased input dimension. These results demonstrate the method’s capability to capture strong non-linearities using appropriately placed knots, even with limited data. Consequently, the method successfully estimated failure probability estimates that are in close agreement with MCS.
However, the current approach may have limited accuracy with uncertain response shapes as it estimates input-output profiles using linear interpolation. In addition, the method cannot currently predict frequency-domain responses in a unified manner. Incorporating higher-order interpolation schemes (e.g., quadratic or cubic) can improve accuracy. Future work will address these limitations by exploring advanced interpolation methods and integrating intrusive dynamical system modifications to enable full-spectrum frequency predictions. Ultimately, the proposed method’s potential in design optimization will be further evaluated by embedding reliability indices within RBDO frameworks.
@@ B.1 Ogden model and Prony series
Ogden model [15] is $U=\sum_{i=1}^{N}\frac{2\delta_{i}}{\alpha_{i}^{2}}\left(\lambda_{1}^{\alpha_{i}}+\lambda_{2}^{\alpha_{i}}+\lambda_{3}^{\alpha_{i}}+\lambda_{4}^{\alpha_{i}}-3\right)+\sum_{i=1}^{N}\frac{1}{D_{i}}\left(U-1\right)^{2},$ where $\bar{\lambda}_{n}$ represents the principal stretches, $J$ is the volume ratio, $\delta_{i}$ and $\alpha_{i}$ are determined by a fitting with the stress-strain curve of the material, and $D_{i}$ means the compressibility. Prony series for viscoelastic [15] is $G(t)=G_{0}\left[1-\sum_{i=1}^{N}g_{i}e^{\left(\frac{-t}{\tau_{i}}\right)}\right],$ where $G(t)$ and $G_{0}$ are the shear modulus and the initial shear modulus. The terms $g_{i}$ and $\tau_{i}$ represent the dimensionless shear-relaxation modulus and the relaxation time.